OpenStudy (anonymous):

FREE MEDAL:::::Let f(x) = x2 - 8x + 5. Find f(-1)

OpenStudy (jdoe0001):

f(-1) just means set x = -1 in \(\bf f(x) = x^2 - 8x + 5\\ f(whatever) = (whatever)^2 - 8(whatever) + 5\)

OpenStudy (anonymous):

what did you get i haven't got an answer?

OpenStudy (jdoe0001):

\(\bf f(x) = x^2 - 8x + 5\\ f(-1) = (-1)^2 - 8(-1) + 5 \implies f(-1) = 1-(-8)+5\)

OpenStudy (jdoe0001):

so, 1+8+5